OpenStudy (anonymous):

Can anyone pleaseeee helppp Let A,B⊆R be nonempty. Define A−B={a−b|a∈A,b∈B}. Prove that inf(A−B)=inf(A)−sup(B

OpenStudy (anonymous):

Have to prove inf(A-B)=inf(A)+inf(-B)

OpenStudy (anonymous):

Then have to show inf(-B)=-sup(B)

OpenStudy (anonymous):

Just not sure how

OpenStudy (abb0t):

@wio

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @cbrezovs Then have to show inf(-B)=-sup(B) \(\color{blue}{\text{End of Quote}}\) If this is true, you'd have to go back to definitions to show it.

OpenStudy (anonymous):

Do u know how to do the problem a diff way?

OpenStudy (anonymous):

I don't remember all the properties of inf and sup.

OpenStudy (anonymous):

inf(A-B)=inf(A)+inf(-B) Do you know this is true, or is it intuitive?

OpenStudy (anonymous):

I believe its true @Zarkon was helping me, he said he knew how to prove it and he told me that was true

OpenStudy (anonymous):

Let's consider the definition. Let \(X\) be defined as \[ \{x| \forall b\in B\quad x<b\} \]That is, all elements below \(B\).\[ x=\inf (B)\iff x\in X\wedge \forall y \in X\quad x\geq y \]

OpenStudy (anonymous):

Here is my attempt at an algebraic definition. It is something we can work with.

OpenStudy (anonymous):

Ok cool

OpenStudy (anonymous):

So we can say \(x_A =\inf (A)\) and \(x_{B^-}=\inf(-B)\)

OpenStudy (anonymous):

Using my definition:\[ \{x| \forall a\in A\quad x<a\} \]\[ x_A=\inf (A)\iff x_A\in X\wedge \forall x \in X\quad x_A\geq x \]And \[ \{x| \forall b\in B\quad x<-b\} \]\[ x_{B^-}=\inf (-B)\iff x_{B^-}\in X\wedge \forall x \in X\quad x_{B^-}\geq x \] And we need to use these facts to show that: \[ \{x| \forall c\in A-B\quad x<c\} \]\[ x_A+x_{B^-}=\inf (A-B)\iff x_A+x_{B^-}\in X\wedge \forall x \in X\quad x_A+x_{B^-}\geq x \]

OpenStudy (anonymous):

We could start by saying that \(c=a-b\)

OpenStudy (anonymous):

Hmmm, I should distinguish those \(X\) sets as well.

OpenStudy (anonymous):

This is a bit messy though, let's start off with something simpler.

OpenStudy (anonymous):

This definition is less messy. We just need to proof stuff about max and lower bounds.\[ X=\{x|\forall a\in A\implies x<a\}\quad x_A=\inf(A)\iff x_A=\max (X) \]

OpenStudy (anonymous):

I think we can show:\[ \max(A-B) =\max(A)+\max(-B) \]By starting with:\[ m_a=\max(A)\iff \forall a\in A\quad m_a\geq a\\ m_b=\max(-B)\iff \forall b'\in -B\quad m_b\geq b' \]First I think we just let \(b'=-b\) and get:\[ \forall b\in B\quad m_b\geq -b \]Then we add these inequalities:\[ \forall a\in A,\;\forall b\in B\quad m_a+m_b\geq a-b \]Whis simplifies to: \[ \forall c\in A-B\quad m_a+m_b\geq c\iff m_a+m_b=\max(A-B) \]

OpenStudy (anonymous):

Is that an m? I'm not sure what that notation is in the last line

OpenStudy (anonymous):

Yes, I use \(m_x\) to be like the max of \(X\) sort of thing.

OpenStudy (anonymous):

I was just trying to show that if you add the max together, you have the max of the sum.

OpenStudy (anonymous):

Let's just get this out of the way \(a\in A,b\in B,a-b=c\in A-B\) I don't like having to type those out all the time.

OpenStudy (anonymous):

Let \( X_A, X_B, X_C \) denote the lower bound of \(A, B, A-B\) respectively.

OpenStudy (anonymous):

You have to show \(x_A=\inf(A),x_B=\inf(-B), x_A+x_B\in X_C\)

OpenStudy (anonymous):

The concept is not hard, but it's just such a tedious proof, unless you are allowed to assume certain properties.

OpenStudy (anonymous):

This is a hint I was g iven use the definition of convergence for each of the sequences, and for convergence of (zn) consider the two cases for n even and odd

OpenStudy (anonymous):

I have to head to bed right now. I appreciate all your help so far.. you've been awesome.So I'm gunna just give u the medal now..maybe we can work on it more tomorrow